Proton Calorimetry/Experimental Runs/2018/Feb21

From PBTWiki

Jump to: navigation, search

Beam tests of the single module with PRaVDA CMOS pixel sensors (PRIAPUS) at the Birmingham Cyclotron with a 36 MeV beam.

Equipment List

Item Notes
Single Module Scintillator Block Optical gel required to couple Scintillator to PMT
Photomultiplier Tube (Hamamatsu 2"?)
Portable Enclosure Modified Peli 1510 Waterproof Wheeled Equipment Case.

230 x 555 x 350mm
Features mount for scintillator and PMT, opening for beam, and ports for SHV, BNC, and SMA cables.
Requires a window (paper? Mylar? tape?) to ensure the opening for the beam is light-tight.

LeCroy HDO6104 Oscilloscope Records PMT output
DAQ Laptop Records HV output / Controls Caen HV supply
Caen NDT1470 HV Supply Supplies HV to PMT
Control Laptop x 2 One each for Remote Desktop control of LeCroy Oscilloscope and DAQ Laptop.
Network Hub x 2 One for the experimental room for the LeCroy HDO6104 and DAQ Laptop to connect to,

one for the control room for each of the Control Laptops to connect to.

Ethernet Cable x 4 To connect each control laptop, the LeCroy HDO6104 and DAQ Laptop to their respective network hubs.
Ethernet Cable To connect the two Network Hubs and complete the network for controlling DAQ equipment from the control room.
SMA (female) to BNC (male) Cable Signal from enclosure port to LeCroy Oscilloscope
SHV Cable HV supply from experimental room to enclosure port
USB Cable Caen HV unit to DAQ Laptop
Gloves For handling scintillator
Optical gel For coupling scintillator to PMT
Wipes For removing optical gel

Beam Tests

Calorimeter files currently located at /unix/pbt/users/dwalker/data/birm_21.02.18 on UCL HEP plus1 server.

PMT Current at HV supply -900 V, beam off: 150 μA

Run Number High Voltage (V) Scope Trigger (self trigger, negative edge, mV) Target Beam Current / Rate Collimator Calorimeter Files Tracker Files Notes
00 -900 -100 ~10 kHz 2 mm run00_2mmCol_160pA_trig100mV --
01 -900 -140 160 pA / ~10 kHz 2 mm run01_2mmCol_160pA_trig140mV 2mmCol_160pA_30s_Run1 Actual beam current rose to 240 pA then 270 pA
02 -900 -70 260 pA 2 mm run02_2mmCol_260pA_trig70mV 2mmCol_260pA_30s_Run2
03 -900 -40 260 pA 2 mm run03_2mmCol_260pA_trig40mV 2mmCol_260pA_30s_Run3 Beam current ~160 pA
04 -900 -140 1 nA 2 mm run04_2mmCol_1nA_trig140mV 2mmCol_1nA_30s_Run4 Beam current ~1.2 nA
Test: HV reduced to -800 V, lights on, no increase in PMT current. Lights off for further beam runs.
05 -900 -140 220 pA 2 mm run05_2mmCol_220pA_trig140mV_1mmPE 2mmCol_220pA_30s_Run5 1 mm polyethylene cover on collimator

Beam current rose to ~260 pA

06 -900 -140 160 pA 2 mm run06_2mmCol_160pA_trig140mV_1mmPE 2mmCol_160pA_30s_Run6 1 mm polyethylene cover on collimator

Beam current rose to ~230 pA

07 -900 -140 10 pA / ~30 kHz 5 mm run07_5mmCol_10pA_trig140mV 5mmCol_10pA_30s_Run7 No polyethylene

Beam current rose to ~30 pA

08 -900 -140 10 pA 5 mm run08_5mmCol_10pA_trig140mV 5mmCol_10pA_30s_Run8 Repeat run for run 07
09 -900 -140 10 pA 5 mm run09_5mmCol_10pA_trig140mV_1mmPE 5mmCol_10pA_30s_Run9 1 mm polyethylene cover on collimator

Beam spot moved - polyethylene cover slipped, 'cyclops' beam Beam current ~8 pA

10 -900 -140 10 pA 5 mm run10_5mmCol_10pA_trig140mV_1mmPE 5mmCol_10pA_30s_Run10 Repeat run for run 09

Peak beam current 15 pA

11 -900 -140 10 pA 5 mm run11_5mmCol_10pA_trig140mV_1mmPE_HalfCover 5mmCol_10pA_30s_Run11 1 mm polyethylene half-cover on collimator

Beam current ~8 pA

Personal tools